OpenStudy (anonymous):

at the art store , marie bought five tubes of oil paint and 3 brushes for $34.75.marcel bought 3 tubes of paint and 1 brush for $17.25 t= tubes b= brush what is the cost of one brush

OpenStudy (anonymous):

i need help plz , porfavor ,s'il vous plaît , per favore whichever language you speak plz help me

OpenStudy (anonymous):

5t + 3b = 34.75 & 3t + b = 17.25, first you need to solve for be by subtracting 3t from 17.25

OpenStudy (anonymous):

Once you have b, you input that into the other equation and find that 5t +3 "b" = 34.75. You can get the solution for a tube of paint. Then you are able to solve for b.

OpenStudy (anonymous):

You can solve either variable first and if you solved for t first. It would eliminate an extra step as you don't have to know how much a tube of paint costs.

OpenStudy (anonymous):

THANK YOU I GAVE U A MEDAL BY THE WAY YOU R SUPER SMART I WISH U CAN BE COUSIN OR MY FRIEND CLOSE FRIEND